OpenStudy (anonymous):

factorize x^3+4x^2+x-6 using the factor theorem. hence solve the eq. x^3+4x^2+x-6=0

OpenStudy (jack1):

ok, so do u know how to factorise these...?

OpenStudy (jack1):

because its a cubic, it will take the form of ( x + __) (x + __) (x + __) ok...?

OpenStudy (jack1):

so what time what times what = -6...? ie what are the factors of 6?

OpenStudy (anonymous):

2*3 & 1*6

OpenStudy (jack1):

yep, perfect

OpenStudy (jack1):

so to "prove"which factors to use, simply sub them in to the equation so start with 2 x^3+4x^2+x-6=0

OpenStudy (jack1):

x^3+4x^2+x-6=0 2^3 + 4(2)^2 + 2 - 6 = 0 8 + 16 + 2 - 6 = 0 20 = 0 so not true, so 2 is not a factor of the eqn however... if we use -2 x^3+4x^2+x-6=0 -8 + 16 + 2 - 6 = 0 0 = 0 so true, so -2 is a factor

OpenStudy (jack1):

so -2 time what times what = -6? must be a 3 and a 1, and one of them must be negative (as answer is negative)

OpenStudy (anonymous):

can i just have the solution of the answer? :D

OpenStudy (jack1):

so lets assume x = +1 sub into equation x^3+4x^2+x-6=0 1 + 4 + 1 - 6 = 0 0 = 0 so true, so 1 is a factor

OpenStudy (jack1):

that means that the other factor must be -3

OpenStudy (jack1):

so try the equation (x - 1) (x + 3) (x + 2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it's right

OpenStudy (jack1):

sweet, there's ur answer then

OpenStudy (anonymous):

ok.. after factoring it.. it asks to solve the eq.